OpenStudy (lovelyharmonics):

Which of the following gases will have the highest velocity at a given temperature?

OpenStudy (lovelyharmonics):

Ne O2 He Cl2

OpenStudy (anonymous):

He ( it has the lowest molar mass) rate Na / rate unknown = square root ( molar mass unknown/ molar mass Ne) 7.0 / 4.9 = square root ( x/ 20) 1.43 = square root ( x / 20) 2.04 = x / 20 x = 41 g/mol

OpenStudy (lovelyharmonics):

lol yeah i know i found that on yahoo.... but i need to know if its right or not

OpenStudy (anonymous):

It's correct!

OpenStudy (lovelyharmonics):

lol how do you know?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

Since you're talking about more than one molecule here, in a given chamber, you would be using an average, and thus, the one with the lightest mass would travel fastest. A bean would jump faster than a rock, if you kicked a table from under.

OpenStudy (abb0t):

Kinetic Energy = mass \(\times\) velotcity

OpenStudy (lovelyharmonics):

lol okay? thanks c: